# An ordering for the strength of functional dependence

### v1

@@ 3 Verifying $\preccurlyeq_{ccx}$
We use the characterizations from the previous section to verify the conditional convex order in various settings. Theorem 1.8 on additive error models is an application of the Schur ordering criterion in Theorem 2.2. In Section 3.1, we use Theorem 1.6 to characterize the conditional convex order for the multivariate normal distribution. In the second part of this section, we apply the dimension reduction principle from Section 2.2 to more general dependence structures.
@@ 3.1 Multivariate normal distribution
The following result characterizes the conditional convex order in the multivariate normal model. Its proof is based on the dimension reduction principle in Example 2.9.
@@
    Due to Theorem 3.1, multivariate normal random vectors $(Y,\mathbb{X})$ and $(Y^{\prime},\mathbb{X}^{\prime})$ are always comparable in conditional convex order, provided $\sigma_{Y}$ and $\sigma_{Y^{\prime}}$ are positive. It is neither assumed that the random vectors have the same dimension nor that the covariance matrices are positive definite. No conditions are imposed on $\bm{\mu}$ and $\bm{\mu}^{\prime}$ either. Due to [2, Proposition 2.7], the conditions in (21) are equivalent to $\xi(Y,\mathbb{X})\leq\xi(Y^{\prime},\mathbb{X}^{\prime})$.
@@
    In Theorem 3.1, the conditional convex order reduces to comparing the Schur complements of the covariance matrices, if $\sigma_{Y}=\sigma_{Y^{\prime}}$. Since $\mathbb{E}[Y\mid\mathbb{X}=\mathbb{x}]=\Sigma_{Y,\mathbb{X}}\Sigma_{\mathbb{X}}^{-}\mathbb{x}$, the conditional convex order in (21) also underlies the fraction of explained variance $\mathrm{Var}(\mathbb{E}[Y\mid\mathbb{X}])/\mathrm{Var}(Y)=\Sigma_{Y,\mathbb{X}}\Sigma_{\mathbb{X}}^{-}\Sigma_{\mathbb{X},Y}/\sigma_{Y}^{2}$ as well as the kernel partial correlation for a linear kernel in [29, Remark 18].
In Example 3.5 below, we verify the ordering condition in (21) for the equicorrelated normal distribution and the multivariate normal distribution with independent predictor variables.
@@ 3.2 Sufficient dimension reduction
The terminology of $g$ being sufficient for $Y|\mathbb{X}$ coincides with the classical concept of a sufficient statistic $g$ for the family of conditional distributions $\{P^{Y|\mathbb{X}=\mathbb{x}}\mid\mathbb{x}\in\mathbb{R}^{p}\}$. For a linear function $g,$ the concept in (22) is also referred to as sufficient for dimension reduction in [18, 41, 42, 40]. An example of a sufficient function for $Y|\mathbb{X}$ is the function $f$ in the additive error model (1) or the linear transformation $f(\mathbb{x})=A\mathbb{x}$ in Example 2.9. Note that (22) is equivalent to conditional independence of $Y$ and $\mathbb{X}$ given $g(\mathbb{X})$; see e.g. [33, Proposition 6.6].
In the following theorem, we provide simple conditions for verifying the conditional convex order under some stochastic monotonicity assumptions. Recall that we write $Y\uparrow_{st}Z$ if $Y$ is SI in $Z$, and similarly, $Y\downarrow_{st}Z$ if $Y$ is stochastically decreasing in $Z$ (i.e., $Y\uparrow_{st}Z$).
@@
The condition $Y\uparrow_{st}g(\mathbb{X})$ is satisfied if there exists a representation $Y=f(g(\mathbb{X}),\varepsilon)$ a.s. for some componentwise increasing function $f$ and an error $\varepsilon$ that is independent of $g(\mathbb{X})$; see e.g. [45, Lemma 3.1]. Without loss of generality, $\varepsilon$ may be uniform on $(0,1)$ or standard normal. The additive error model (1) admits such a representation. Note that the proof of Theorem 1.8 applies Theorem 3.3(i) but not the criteria on pointwise comparison of copulas in parts (ii) and (iii) because the distributions of $g(\mathbb{X})$ and $\varepsilon$ in the additive error model are not given and thus the copula $C_{Y,g(\mathbb{X})}$ is not known.
In the next example, we use Theorem 3.3 to verify the $\preccurlyeq_{ccx}$-ordering criterion in (21) for two specific families of multivariate normal distributions.
@@
    The copula of $(Y,\sum X_{i})$ is the Gaussian copula with correlation $\varrho\,.$ Since the Gaussian copula family is $\leq_{c}$-increasing in its parameter and since $Y\uparrow_{st}\sum X_{i}$ for $\varrho\geq 0$, we obtain from Theorem 3.3(ii) that $(Y,\mathbb{X})$ is increasing in $\varrho$ with respect to the conditional convex order. Using radial symmetry of the bivariate normal distribution (i.e., $(-Y,-\sum X_{i})\stackrel{{\scriptstyle\mathrm{d}}}{{=}}(Y,\sum X_{i})$), using $Y\downarrow_{st}\sum X_{i}$ for $\varrho\leq 0$, and applying Theorem 3.3(iii), we obtain that $(Y,\mathbb{X})$ is increasing eventually in $|\varrho|$ with respect to $\preccurlyeq_{ccx}$. Note that, by (21), this means that $\Sigma_{Y,\mathbb{X}}\Sigma_{\mathbb{X}}^{-}\Sigma_{\mathbb{X},Y}$ is increasing in $|\varrho|$.
@@
    Consider the case where the components of $\mathbb{X}=(X_{1},\ldots,X_{p})$ are independent and where $\mathrm{Cor}(Y,X_{i})=\varrho$ for all $i$. Then the covariance matrix $\Sigma$ is positive semi-definite if and only if $\varrho\in[-p^{-1/2},p^{-1/2}]$. Again, the conditional distribution
@@
    Similar as above, we obtain from Theorem 3.3(ii),(iii) that $(Y,\mathbb{X})$ is increasing in $|\varrho|$ with respect to the conditional convex order.
In the following example, we outline various well known SI copula families that are increasing in the concordance order. This yields families of models $Y=f(g(\mathbb{X}),\varepsilon)$ that fulfill the assumptions of Theorem 3.3(ii) and thus are comparable in conditional convex order; see Remark 3.4. By some symmetry arguments, similar results can be obtained for stochastically decreasing copulas.
@@
Let $Y=f(g(\mathbb{X}),\varepsilon)$ with copula $C_{1}=C_{Y,g(\mathbb{X})}$ and $Y^{\prime}=f(g(\mathbb{X}^{\prime}),\varepsilon^{\prime})$ with copula $C_{2}=C_{Y^{\prime},g(\mathbb{X}^{\prime})}$. Assume that $\overline{\mathsf{Ran}(F_{Y})}=\overline{\mathsf{Ran}(F_{Y^{\prime}})}$. Then, we can verify $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$ by Theorem 3.3(ii) for Archimedean, extreme-value, and elliptical copulas as follows.
@@
    Assume that $C_{i}$ is an Archimedean copula, i.e., $C_{i}$ admits a representation $C_{i}(u,v)=\psi_{i}(\varphi_{i}(u),\varphi_{i}(v))$ for some decreasing convex function $\varphi_{i}\colon[0,1]\to[0,\infty)$ with $\varphi(1)=0$, where $\psi_{i}=\varphi_{i}^{[-1]}$ denotes the (pseudo-)inverse of $\varphi_{i}$. If $\psi_{i}^{\prime}$ is log-convex, then $C_{i}$ is SI ([46, Theorem 2.8]; further, if $\varphi_{1}\circ\psi_{2}$ is subadditive, then $C_{1}\leq_{c}C_{2}$ ([48, Theorem 4.4.2]). Hence, under these conditions, we have $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$. Various Archimedean copula families that meet these conditions are given in [3, Table 3].
@@
    Assume that $C_{i}$ is an extreme-value copula, i.e., $C_{i}$ admits a representation
    $C_{i}(u,v)=(\exp(\log(uv))A_{i}(\log(v)/(\log(u)+\log(v))$ where the Pickands dependence function $A_{i}$ is convex and satisfies the constraints $\max\{t,1-t\}\leq A_{i}(t)\leq 1$ for all $t\in[0,1]$. Extreme-value copulas are always SI; moreover, $A_{1}(t)\geq A_{2}(t)$ for all $t\in(0,1)$ implies $C_{1}\leq_{c}C_{2}$; see [3, Theorem 3.4]. Various well known extreme-value copula families that meet the conditions are given in [3, Table 4].
@@
    Assume that $C_{i}$ is an elliptical copula, i.e., it is the copula (implicitly obtained by Sklar’s theorem) of an elliptically distributed bivariate random vector $\mathbb{Z}_{i}$, that is, $\mathbb{Z}_{i}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}R_{i}A_{i}\mathbb{U}$ for a non-negative random variable $R_{i}$ that is independent of the bivariate random vector $\mathbb{U}$ which is uniformly distributed on the $2$-sphere. Here, $A_{i}$ is a $2\times 2$-matrix. Assume that $\Sigma_{i}:=A_{i}A_{i}^{T}=\begin{pmatrix}1&\rho_{i}\\ \rho_{i}&1\end{pmatrix}$ for $\rho_{i}\in[-1,1]$. Sufficient conditions on $R_{i}$ and $\rho_{i}$ for $C_{i}$ being SI are given in [1, Proposition 1.2]. Further, if $\rho_{1}\leq\rho_{2}$ then $C_{1}\leq_{c}C_{2}$. Elliptical SI copula families that meet these conditions are given in [3, Table 5].
@@ 4 Dependence measures generated by $\preccurlyeq_{ccx}$
In this section, we use the conditional convex order to generate new dependence measures. While convexity guarantees that information monotonicity is inherited from $\leq_{ccx}$ to the corresponding dependence measure, the zero-independence and max-functionality properties require additional regularity conditions; compare [5, 13]. We further provide simple conditions under which these dependence measures characterize conditional independence.
@@
$\Lambda_{\varphi}(Y,\mathbb{X}):$ $=\beta_{\varphi}^{-1}\int_{\mathbb{R}^{2p}}\int_{\mathbb{R}}\varphi(F_{Y|\mathbb{X}=\mathbb{x}_{1}}(y)-F_{Y|\mathbb{X}=\mathbb{x}_{2}}(y))\mathsf{\,d}P^{Y}(y)\mathsf{\,d}(P^{\mathbb{X}}\otimes P^{\mathbb{X}})(\mathbb{x}_{1},\mathbb{x}_{2})$
with normalizing constants $\alpha_{\varphi}:=\int_{\mathbb{R}}\int_{\mathbb{R}}\varphi\left(\mathds{1}_{\{y_{1}\leq y\}}-F_{Y}(y)\right)\mathsf{\,d}P^{Y}(y)\mathsf{\,d}P^{Y}(y_{1})$ and
$\beta_{\varphi}:=\int_{\mathbb{R}^{2}}\int_{\mathbb{R}}\varphi\left(\mathds{1}_{\{y_{1}\leq y\}}-\mathds{1}_{\{y_{2}\leq y\}}\right)\mathsf{\,d}P^{Y}(y)\mathsf{\,d}(P^{Y}\otimes P^{Y})(y_{1},y_{2})\,.$ While $\xi_{\varphi}$ compares conditional distribution functions with unconditional ones, $\Lambda_{\varphi}$ measures the sensitivity of conditional distribution functions in the conditioning variable.
By the following theorem, which is based on the Hardy–Littlewood–Polya theorem (see Lemma 2.1), convexity conditions on $\varphi$ imply that $\xi_{\varphi}$ and $\Lambda_{\varphi}$ are dependence measure that are $\preccurlyeq_{ccx}$-increasing and characterize (conditional) independence and perfect dependence. Note that, for $\varphi(x)=x^{2}$, the functionals $\xi_{\varphi}$ and $\Lambda_{\varphi}$ reduce to Chatterjee’s rank correlation $\xi$ defined in Equation (2).
@@
Assume that $\varphi:[-1,1]\rightarrow\mathbb{R}$ is convex and strictly convex in 0 with $\varphi(0)=0$. Assume that $Y$ is non-degenerate. Then $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$ implies
@@
Under the assumptions of Theorem 4.1,
@@
    The normalizing constants $\alpha_{\varphi}$ and $\beta_{\varphi}$ are positive. Note that the marginal constraint implies that also $Y^{\prime}$ is non-degenerate.
@@
    $\xi_{\varphi}$ and $\Lambda_{\varphi}$ are dependence measures that inherit information monotonicity and invariance properties from the conditional convex order, i.e., $\xi_{\varphi}(Y,\mathbb{X})\leq\xi_{\varphi}(Y,(\mathbb{X},\mathbb{Z}))$, $\xi_{\varphi}(Y,\mathbb{X})=\xi_{\varphi}(g(Y),\mathbf{h}(\mathbb{X}))$ for all strictly increasing functions $g$ and bijective functions $\mathbf{h}$, and $\xi_{\varphi}(Y,\mathbb{X})=\xi_{\varphi}(F_{Y}(Y),\mathbf{F}_{\mathbb{X}}(\mathbb{X}))$, similarly for $\Lambda_{\varphi}$;
@@
    Strict convexity of $\varphi$ in $\varphi(0)=0$ is sufficient for the values 0 and $1$ to characterize independence and perfect dependence, respectively. However, strict convexity in 0 is not sufficient for $\xi_{\varphi}$ to characterize conditional independence. As a counterexample, consider $\varphi\colon t\mapsto|t|$ which is convex and strictly convex in $\varphi(0)=0$. In this case, $\xi_{\varphi}(Y,\mathbb{X})=\xi_{\varphi}(Y,(\mathbb{X},\mathbb{Z}))$ does not imply conditional independence of $Y$ and $\mathbb{Z}$ given $\mathbb{X}$; see [27, Example 5.10].
@@
    The zero-independence and max-functionality property as well as information monotonicity for $\xi_{\varphi}$ can be traced back to convexity properties for measures of statistical association as studied in [13]. In Theorem 4.1, we also provide a characterization of conditional independence and show that $\xi_{\varphi}$ is increasing in $\preccurlyeq_{ccx}$.

### v2

@@ 3 Verifying $\preccurlyeq_{ccx}$
In this section, we provide sufficient conditions for verifying the conditional convex order. These conditions are based on the dimension reduction principle from Section 2.2 and on stochastic monotonicity assumptions which simplify a verification of the Schur order criterion in Theorem 2.2. In the first part of this section, we characterize the $\preccurlyeq_{ccx}$-order for multivariate normal distributions. In the second part, we verify the conditional convex order for various copula-based models that generalize the additive error model in (1).
@@ 3.1 Multivariate normal distribution
The following result characterizes $\preccurlyeq_{ccx}$ for the multivariate normal distribution. Its proof is based on the dimension reduction performed in Example 2.9.
@@
    Due to Theorem 3.1, multivariate normal random vectors $(Y,\mathbb{X})$ and $(Y^{\prime},\mathbb{X}^{\prime})$ are always comparable in $\preccurlyeq_{ccx}$, provided $\sigma_{Y}$ and $\sigma_{Y^{\prime}}$ are positive. It is neither assumed that the random vectors have the same dimension nor that the covariance matrices are positive definite. No conditions are imposed on $\bm{\mu}$ and $\bm{\mu}^{\prime}$ either. Interestingly, by [3, Proposition 2.7], either condition in (24) is equivalent to $\xi(Y,\mathbb{X})\leq\xi(Y^{\prime},\mathbb{X}^{\prime})$.
@@
    In Theorem 3.1, if $\sigma_{Y}=\sigma_{Y^{\prime}}$, the conditional convex order reduces to comparing the Schur complements of the covariance matrices. Since $\mathbb{E}[Y\mid\mathbb{X}=\mathbb{x}]=\Sigma_{Y,\mathbb{X}}\Sigma_{\mathbb{X}}^{-}\mathbb{x}$, $\preccurlyeq_{ccx}$ in (24) also underlies the fraction of explained variance $\mathrm{Var}(\mathbb{E}[Y\mid\mathbb{X}])/\mathrm{Var}(Y)=\Sigma_{Y,\mathbb{X}}\Sigma_{\mathbb{X}}^{-}\Sigma_{\mathbb{X},Y}/\sigma_{Y}^{2}$ and the kernel partial correlation for a linear kernel [29, Remark 18].
In Example 3.5 below, we verify the ordering condition (24) for the equicorrelated normal distribution and the multivariate normal distribution with independent predictor variables.
@@ 3.2 Sufficient dimension reduction
The terminology of $g$ being sufficient for $Y|\mathbb{X}$ coincides with the classical concept of a sufficient statistic $g$ for the family of conditional distributions $\{P^{Y|\mathbb{X}=\mathbb{x}}\mid\mathbb{x}\in\mathbb{R}^{p}\}$. For a linear function $g,$ the concept in (25) is also referred to as sufficient for dimension reduction in [18, 40, 41, 39]. An example of a sufficient function for $Y|\mathbb{X}$ is the function $f$ in the additive error model (1) or the linear transformation $g(\mathbb{x})=A\mathbb{x}$ in Example 2.9. Note that (25) is equivalent to conditional independence of $Y$ and $\mathbb{X}$ given $g(\mathbb{X})$; see [33, Proposition 6.6].
In the following theorem, we provide simple conditions for verifying $\preccurlyeq_{ccx}$ under stochastic monotonicity assumptions. We write $Y\uparrow_{st}Z$ if $Y$ is SI in $Z$, and similarly, $Y\downarrow_{st}Z$ if $Y$ is stochastically decreasing in $Z$ (i.e., $Y\uparrow_{st}Z$).
@@
  1. (a)
    The condition $Y\uparrow_{st}g(\mathbb{X})$ is satisfied if there exists a representation $Y=f(g(\mathbb{X}),\varepsilon)$ a.s. for some componentwise increasing function $f$ and an error $\varepsilon$ that is independent of $g(\mathbb{X})$; see e.g. [46, Lemma 3.1]. Without loss of generality, $\varepsilon$ may be uniform on $(0,1)$ or standard normal. The additive error model (1) admits such a representation. Note that the proof of Theorem 1.8 applies Theorem 3.3i but not the criteria on pointwise comparison of copulas in parts ii and iii because the distributions of $g(\mathbb{X})$ and $\varepsilon$ in the additive error model are not given and thus the copula $C_{Y,g(\mathbb{X})}$ is not known.
  2. (b)
    The stochastic monotonicity assumptions in Theorem 3.3 facilitate the verification of the conditional convex order because, for bivariate SI random vectors from the same Fréchet class, $\preccurlyeq_{ccx}$ and $\leq_{c}$ are equivalent; see Proposition 5.1. For the latter concordance order, numerous comparison results are available for parametric families of distributions; see Example 3.6 below. Note that Proposition 2.5 on Bernoulli distributions does not require stochastic monotonicity assumptions as in Theorem 3.3.
In the next example, we use Theorem 3.3 to verify for two specific families of multivariate normal distributions the $\preccurlyeq_{ccx}$-ordering criterion in (24).
@@
    The copula of $(Y,\sum X_{i})$ is the Gaussian copula with correlation $\varrho\,.$ Since the Gaussian copula family is $\leq_{c}$-increasing in its parameter and since $Y\uparrow_{st}\sum X_{i}$ for $\varrho\geq 0$, we obtain from Theorem 3.3ii that $(Y,\mathbb{X})$ is increasing in $\varrho$ with respect to the conditional convex order. Using radial symmetry of the bivariate normal distribution (i.e., $(-Y,-\sum X_{i})\stackrel{{\scriptstyle\mathrm{d}}}{{=}}(Y,\sum X_{i})$), using $Y\downarrow_{st}\sum X_{i}$ for $\varrho\leq 0$, and applying Theorem 3.3iii, we obtain that $(Y,\mathbb{X})$ is increasing eventually in $|\varrho|$ with respect to $\preccurlyeq_{ccx}$. Note that, by (24), this means that $\Sigma_{Y,\mathbb{X}}\Sigma_{\mathbb{X}}^{-}\Sigma_{\mathbb{X},Y}$ is increasing in $|\varrho|$.
@@
    Consider the case where $\mathrm{Var}(Y)=1$, the components of $\mathbb{X}=(X_{1},\ldots,X_{p})$ are independent, and $\mathrm{Cor}(Y,X_{i})=\varrho$ for all $i$. Then the covariance matrix $\Sigma$ is positive semi-definite if and only if $\varrho\in[-p^{-1/2},p^{-1/2}]$. Again, the conditional distribution
@@
    Similar as above, we obtain from Theorem 3.3ii and iii that $(Y,\mathbb{X})$ as well as the Schur complements $\Sigma_{Y,\mathbb{X}}\Sigma_{\mathbb{X}}^{-}\Sigma_{\mathbb{X},Y}$ are increasing in $|\varrho|$ with respect to $\preccurlyeq_{ccx}$.
In the following example, we outline various well known SI copula families that are increasing in the concordance order. This yields families of models $Y=f(g(\mathbb{X}),\varepsilon)$ that fulfill the assumptions of Theorem 3.3ii and thus are $\preccurlyeq_{ccx}$-comparable; see Remark 3.4. Symmetry arguments yield similar results for stochastically decreasing copulas.
@@
Let $Y=f(g(\mathbb{X}),\varepsilon)$ with copula $C_{1}=C_{Y,g(\mathbb{X})}$ and $Y^{\prime}=f(g(\mathbb{X}^{\prime}),\varepsilon^{\prime})$ with copula $C_{2}=C_{Y^{\prime},g(\mathbb{X}^{\prime})}$. Assume that $\overline{\mathsf{Ran}(F_{Y})}=\overline{\mathsf{Ran}(F_{Y^{\prime}})}$. Then, we can verify $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$ by Theorem 3.3ii for Archimedean, extreme-value, and elliptical copulas $C_{1}$ and $C_{2}$ as follows.
@@
    Assume that $C_{i}$ is an Archimedean copula, i.e., $C_{i}$ admits a representation $C_{i}(u,v)=\psi_{i}(\varphi_{i}(u),\varphi_{i}(v))$ for some decreasing convex function $\varphi_{i}\colon[0,1]\to[0,\infty)$ with $\varphi(1)=0$, where $\psi_{i}=\varphi_{i}^{[-1]}$ denotes the (pseudo-)inverse of $\varphi_{i}$. If $\psi_{i}^{\prime}$ is log-convex, then $C_{i}$ is SI [47, Theorem 2.8]; further, if $\varphi_{1}\circ\psi_{2}$ is subadditive, then $C_{1}\leq_{c}C_{2}$ [48, Theorem 4.4.2]. Various Archimedean copula families that meet these conditions are given in [5, Table 3].
@@
    Assume that $C_{i}$ is an extreme-value copula, i.e., $C_{i}$ admits a representation $C_{i}(u,v)=(\exp(\log(uv))A_{i}(\log(v)/(\log(u)+\log(v))$ where the Pickands dependence function $A_{i}$ is convex and satisfies the constraints $\max\{t,1-t\}\leq A_{i}(t)\leq 1$ for all $t\in[0,1]$. Extreme-value copulas are always SI; moreover, $A_{1}(t)\geq A_{2}(t)$ for all $t\in(0,1)$ implies $C_{1}\leq_{c}C_{2}$; see [5, Theorem 3.4]. Various well known extreme-value copula families that meet the conditions are given in [5, Table 4].
@@
    Assume that $C_{i}$ is an elliptical copula, i.e., it is the copula (implicitly obtained by Sklar’s theorem) of an elliptically distributed bivariate random vector $\mathbb{Z}_{i}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}RA_{i}\mathbb{U}$ for a non-negative random variable $R$ that is independent of the bivariate random vector $\mathbb{U}$ which is uniformly distributed on the $2$-sphere. Here, $A_{i}$ is a $2\times 2$-matrix such that $\Sigma_{i}:=A_{i}A_{i}^{T}=\begin{pmatrix}1&\rho_{i}\\ \rho_{i}&1\end{pmatrix}$ for $\rho_{i}\in[-1,1]$. Sufficient conditions on $R_{i}$ and $\rho_{i}$ for $C_{i}$ being SI are given in [1, Proposition 1.2]. Further, if $\rho_{1}\leq\rho_{2}$ then $C_{1}\leq_{c}C_{2}$. Elliptical SI copula families that meet these conditions are given in [5, Table 5].
@@ 4 Dependence measures generated by $\preccurlyeq_{ccx}$
In this section, we use the conditional convex order to construct dependence measures based on conditional distribution functions, Wasserstein distances, and monotone rearrangements. These three classes correspond to the three equivalences in Theorem 1.6. The resulting dependence measures inherit information monotonicity and invariance properties from $\preccurlyeq_{ccx}$. The zero-independence and max-functionality property require some regularity conditions based on convexity; see [4, 12]. We further provide simple conditions under which these dependence measures characterize conditional independence. As noted in Remark 3.2b, $\preccurlyeq_{ccx}$ also underlies the kernel partial correlation [29] for a linear kernel.
Throughout, we assume $Y$ to be non-degenerate so that the proposed dependence measures attain at least two distinct values; if $Y$ were degenerate, it would be both independent of $\mathbb{X}$ and perfectly dependent on $\mathbb{X}$, contradicting zero-independence and max-functionality.
4.1 Dependence measures based on conditional distribution functions
@@
$\Lambda_{\varphi}(Y,\mathbb{X}):$ $=\beta_{\varphi}^{-1}\int_{\mathbb{R}^{2p}}\int_{\mathbb{R}}\varphi(F_{Y|\mathbb{X}=\mathbb{x}_{1}}(y)-F_{Y|\mathbb{X}=\mathbb{x}_{2}}(y))\mathsf{\,d}P^{Y}(y)\mathsf{\,d}(P^{\mathbb{X}}\otimes P^{\mathbb{X}})(\mathbb{x}_{1},\mathbb{x}_{2})$ (28)
with normalizing constants $\alpha_{\varphi}:=\int_{\mathbb{R}}\int_{\mathbb{R}}\varphi\left(\mathds{1}_{\{y_{1}\leq y\}}-F_{Y}(y)\right)\mathsf{\,d}P^{Y}(y)\mathsf{\,d}P^{Y}(y_{1})$ and $\beta_{\varphi}:=\int_{\mathbb{R}^{2}}\int_{\mathbb{R}}\varphi\left(\mathds{1}_{\{y_{1}\leq y\}}-\mathds{1}_{\{y_{2}\leq y\}}\right)\mathsf{\,d}P^{Y}(y)\mathsf{\,d}(P^{Y}\otimes P^{Y})(y_{1},y_{2})\,.$ While $\xi_{\varphi}$ compares conditional distribution functions with unconditional ones, $\Lambda_{\varphi}$ measures the sensitivity of conditional distribution functions in the conditioning variable. Note that, for $\varphi(x)=x^{2}$, the functionals $\xi_{\varphi}$ and $\Lambda_{\varphi}$ reduce to Chatterjee’s rank correlation $\xi$ defined in (2).
By the following result, which is based on the Hardy–Littlewood–Polya theorem (see Lemma 2.1), convexity conditions on $\varphi$ imply that $\xi_{\varphi}$ and $\Lambda_{\varphi}$ are dependence measure that are $\preccurlyeq_{ccx}$-increasing and characterize (conditional) independence and perfect dependence.
@@
Assume that $\varphi:[-1,1]\rightarrow\mathbb{R}$ is convex and strictly convex at 0 with $\varphi(0)=0$. Then $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$ implies
@@
    $\xi_{\varphi}$ and $\Lambda_{\varphi}$ satisfy information monotonicity, i.e. $\xi_{\varphi}(Y,\mathbb{X})\leq\xi_{\varphi}(Y,(\mathbb{X},\mathbb{Z}))$ and $\Lambda_{\varphi}(Y,\mathbb{X})\leq\Lambda_{\varphi}(Y,(\mathbb{X},\mathbb{Z}))$.
  4. (iv)
@@
Suppose the assumptions of Theorem 4.1.
@@
    The normalizing constants $\alpha_{\varphi}$ and $\beta_{\varphi}$ in (28) are positive. Note that the marginal constraint implies that also $Y^{\prime}$ is non-degenerate.
@@
    $\xi_{\varphi}$ and $\Lambda_{\varphi}$ are dependence measures that inherit all invariance properties from $\preccurlyeq_{ccx}$, i.e. $\xi_{\varphi}(Y,\mathbb{X})=\xi_{\varphi}(g(Y),\mathbf{h}(\mathbb{X}))$ for all strictly increasing functions $g$ and bijective functions $\mathbf{h}$, and $\xi_{\varphi}(Y,\mathbb{X})=\xi_{\varphi}(F_{Y}(Y),\mathbf{F}_{\mathbb{X}}(\mathbb{X}))$, similarly for $\Lambda_{\varphi}$.
@@
    Strict convexity of $\varphi$ at $\varphi(0)=0$ is sufficient for the values 0 and $1$ to characterize independence and perfect dependence, respectively. However, strict convexity at 0 is not sufficient for $\xi_{\varphi}$ to characterize conditional independence. As a counterexample, consider $\varphi\colon t\mapsto|t|$ which is convex and strictly convex at $\varphi(0)=0$. In this case, $\xi_{\varphi}(Y,\mathbb{X})=\xi_{\varphi}(Y,(\mathbb{X},\mathbb{Z}))$ does not imply conditional independence of $Y$ and $\mathbb{Z}$ given $\mathbb{X}$; see [27, Example 5.10].
@@
    The zero-independence and max-functionality property as well as information monotonicity for $\xi_{\varphi}$ can be traced back to convexity properties for measures of statistical association as studied in [12]. In Theorem 4.1, we also provide a characterization of conditional independence and show that $\xi_{\varphi}$ is $\preccurlyeq_{ccx}$-increasing.